### v1

# Measuring Ransomware Lateral Movement Susceptibility via Privilege-Weighted Adjacency Matrix Exponentiation Satyam Tyagi1 Ganesh Murugesan1

Abstract: Ransomware impact hinges on how easily an intruder can move laterally and spread to the maximum number of assets. We present a graph-theoretic method to measure lateral-movement susceptibility and estimate blast radius. We build a directed multigraph where vertices represent assets and edges represent reachable services (e.g., RDP/SSH) between them. We model lateral movement as a probabilistic process using a pivot potential factor $\pi(s)$ for each service. This allows us to iteratively compute a $K$-hop compromise probability matrix that captures how compromise propagates through the network. Metrics derived from this model include: (1) Lateral-Movement Susceptibility (LMSK): the average probability of a successful lateral movement between any two assets (0-1 scale); and (2) Blast-Radius Estimate (BREK): the expected percentage of assets compromised in an average attack scenario. Interactive control (SSH 22, RDP 3389) gets higher $\pi(s)$ than app-only ports (MySQL 3306, MSSQL 1433), which seldom enable pivoting without an RCE. Across anonymized enterprise snapshots, pruning high-$\pi(s)$ edges yields the largest LMSK/BREK drop, aligning with CISA guidance, MITRE ATT&CK (TA0008: Lateral Movement), and NIST SP 800-207. The framework evaluates (micro)segmentation and helps prioritize controls that reduce lateral movement susceptibility and shrink blast radius.
Keywords: adjacency matrix; multi-hop reachability; privilege weighting; lateral movement; microsegmentation; blast radius; ransomware; Zero Trust
Mathematics Subject Classification : 05C50; 05C90; 94C15
1.  Introduction
Ransomware impact is determined less by where an adversary first lands and more by how easily they can move laterally to reach high-value assets and propagate encryption. Classic reachability (Boolean adjacency, $k$-hop closure) reveals where paths exist, but it can overstate risk by treating an SSH or RDP foothold as equivalent to a low-privilege application socket. In practice, interactive remote services (e.g., RDP/SSH) enable immediate pivoting; app-only ports typically demand additional exploits. Our goal is a probability-weighted alternative that measures the ease of movement—and the expected blast radius—using only topology plus coarse service-class information.
We introduce a privilege- and account-aware adjacency model that composes per-service “pivot potentials” $\pi(s)\!\in\![0,1]$ with multiplicative step success and a noisy-OR union across alternative routes. Iterating this algebraic path closure yields (i) a $k$-hop (or $\varepsilon$-fixed-point) compromise probability matrix and (ii) two deployment metrics: LMSK (lateral-movement susceptibility; mean off-diagonal) and BREK (blast-radius estimate; mean including the diagonal).
CVE-free lateral movement via valid accounts.
Once an adversary holds valid credentials and an interactive remote service is exposed (RDP/SSH), lateral movement does not require unpatched vulnerabilities: adversaries log into remote services and act as the authenticated user [MITRE, 2025f, b, c]. Patch status is orthogonal to pivotability when privileged landing is possible; hence our operators weight service exposure and privilege class, not CVE likelihoods.
Contributions.
  • Privilege- and account-aware path closure. We formalize lateral movement over $[0,1]$ with step-wise multiplication and noisy-OR unions, explicitly capturing valid-account pivots on interactive services [MITRE, 2025b].
  • Service-class lifting. We encode exact “land-then-pivot” semantics by lifting states to $(\text{node},\text{class})$, preserving first-hop reach while applying class-conditioned pivot probabilities thereafter.
  • Actionable metrics. We define LMSK and BREK for policy comparison and segmentation planning.
  • Theoretical bounds. We situate our iteration within algebraic path theory, giving a series upper bound via $U_{K}=A+AW+\cdots+AW^{K-1}$ and a single-path product lower bound.
  • Alignment with guidance. Our empirical reductions in LMSK/BREK under RDP/SSH constraints match Zero Trust and ransomware guidance to limit remote desktop services and constrain lateral movement [Rose et al., 2020; CISA, 2025a, b, c].
Relation to Zero Trust and segmentation.
Zero Trust (NIST SP 800-207) calls for minimizing implicit trust and continuously evaluating access in context [Rose et al., 2020]. Our metrics quantify the residual ease of movement under candidate policies, complementing microsegmentation by highlighting edges whose removal yields the steepest probability-mass drop.
Organization.
Section 2 reviews the literature. Section 3 formalizes matrices, operators, and fixed-point computation. Section 4 defines LMSK/BREK. Section 5 presents properties and bounds. Section 6 provides examples and evidence.
2.  Background and Related Work
Account-based pivots vs. vulnerability-driven traversal. Beyond exploit-based movement (e.g., T1210) [MITRE, 2025d], ATT&CK catalogs techniques enabling lateral movement without software flaws: Valid Accounts (T1078) and Remote Services (T1021, including RDP/SSH) describe abusing legitimate credentials to authenticate and operate on remote hosts [MITRE, 2025f, b, c, e]. Our formulation treats interactive services as high-pivot edges when paired with account access, independent of CVE posture.
Attack graphs and probabilistic security metrics. Classical attack graphs model preconditions/exploits and enumerate paths to goals [Phillips and Swiler, 1998; Sheyner et al., 2002]. Later work augments graphs with probabilities and costs. We deliberately do not encode per-CVE logic; instead we treat reachable services as coarse pivot opportunities whose probabilities compose multiplicatively with noisy-OR unions, trading granularity for deployability on routine network snapshots.
Algebraic path problems and semiring views. Many path computations admit uniform treatment via algebraic path frameworks replacing $(+,\times)$ by problem-specific $(\oplus,\otimes)$. We adopt multiplication for per-step success and probabilistic sum (noisy-OR) for unions, yielding a monotone iteration toward a least fixed point; this aligns with semiring-style closures, though our $\oplus$ is non-idempotent [Mohri, 2002; Sanmartín and Naranjo, 2022].
Noisy-OR and causal independence. Our union operator is the standard noisy-OR from probabilistic graphical models, aggregating independent causal mechanisms with low elicitation burden [Heckerman, 1994]. Generalizations (e.g., Noisy-MAX) clarify when such factorizations are appropriate.
Epidemic spreading analogies. Multiplicative transmission along edges with union across parallel paths is mathematically analogous to network contagion; classic results connect topology to spread [Newman, 2002]. Our aim is risk measurement (expected reach/blast), not epidemic prediction, but the analogy explains the outsized effect of a small set of high-pivot edges.
AD attack-path tooling. In enterprise Windows domains, BloodHound operationalizes graph-based lateral-movement analysis at scale [SpecterOps, 2025]. We follow similar instincts (directed, typed edges) but compute a probability-weighted closure and provide system-level metrics for control evaluation.
Zero Trust and ransomware guidance. NIST SP 800-207 sets the architectural principle; ransomware advisories emphasize limiting RDP/remote services to reduce both initial access and lateral movement [Rose et al., 2020; CISA, 2025a, b, c].
2.1 Graph-theoretic underpinnings
$k$-hop walks via adjacency powers.
Let $A$ be the (directed) adjacency matrix of the network. Classical algebraic graph theory shows that $(A^{k})_{ij}$ equals the number of $i{\to}j$ walks of length kk. [^1] As usual, walks may repeat vertices; paths are walks with no repeated vertices. Our probabilistic closure works with walks. See, e.g., Godsil and Royle [2001]; Biggs [1993]; Diestel [2005]. This formalizes the “$k$-hop” notion we use when bounding series $A+AW+\cdots+AW^{K-1}$.
Boolean-OR reachability and transitive closure.
Over the Boolean semiring $(\lor,\land)$, the transitive closure (reachability) can be computed by Boolean matrix powering and union,
$R\;=\;A\lor A^{(2)}\lor\cdots\lor A^{(n)}\,$,
or equivalently via Warshall’s dynamic program for transitive closure [Warshall, 1962]. From a more algebraic perspective, Kleene-style closure over idempotent semirings and dioids gives a unifying view of reachability and shortest paths [Kozen, 1994; Gondran and Minoux, 2008].
Noisy-OR style union for probabilistic graphs.
When parallel, conditionally independent mechanisms can each enable a step (e.g., multiple services/credentials offering a pivot), the probability that at least one succeeds is the probabilistic sum
$p_{\text{union}}\;=\;1-\prod_{\ell}(1-p_{\ell})\,$,
the same aggregation used in network reliability for parallel components [Colbourn, 1987]. In probabilistic graphical models, this is the standard noisy-OR (and its generalizations), which captures causal independence with one parameter per parent [Pearl, 1988; Heckerman, 1994; Diez, 1993]. Our use of a noisy-OR union along with multiplicative step success places the model within these classical frameworks.
3.  The Proposed Model
Our model quantifies lateral movement exposure by representing the network as a directed graph and simulating the propagation of compromise probability. This requires two foundational matrices derived from network state and security policies.
3.1 Foundational Matrices
  1. 1.
    The Adjacency Matrix ($A$): This is an $n\times n$ binary matrix representing the network’s underlying structure, where $n$ is the number of assets. An edge exists only if a service is both permitted by network policy and actively listening on the destination asset.
[…]
    We take $A_{ii}=0$ for all $i$ (no self-loops). Let $I$ denote the $n\times n$ identity matrix.
  2. 2.
    The Pivot Potential Matrix ($W$): This is an $n\times n$ matrix that encodes the inherent pivot potential of each edge. Instead of considering only the single most powerful service, the model treats each reachable service on an edge as an independent opportunity for an attacker. The entry $(W)_{ij}$ represents the total probability of a successful pivot from asset $i$ to $j$, calculated as the probabilistic union (or ”noisy-OR”) of the pivot potentials of all services available on that edge.
    Let $S_{ij}$ be the set of all services reachable from asset $i$ to asset $j$. The pivot potential is calculated as:
    $W_{ij}=1-\prod_{s\in S_{ij}}(1-\pi(s))$
    Convention. $S_{ij}$ includes only services that are (i) allowed by policy and (ii) observed listening on $j$; hence $W_{ij}>0\Rightarrow A_{ij}=1$ (i.e., $W\leq A$ elementwise). In particular, if $S_{ij}=\varnothing$ then $W_{ij}=0$.
3.2 Computing Reachability (Unweighted Method)
Before detailing our probabilistic model, we first describe the standard, unweighted method for calculating K-hop reachability. This method answers the binary question: ”Does a path of length $K$ or less exist?” This is computed iteratively. Let $S_{k}$ be the matrix where $(S_{k})_{ij}=1$ if a path of at most $k$ hops exists.
  • Base Case: The 1-hop reachability matrix is the adjacency matrix, $S_{1}=A$.
  • Iteration: The reachability matrix for at most $k$ hops, $S_{k}$, is found by taking the previous state $S_{k-1}$ and adding any new paths formed by one additional hop:
    $S_{k}=S_{k-1}\lor(S_{k-1}\cdot_{bool}A)$
  • Final Result: The result is the matrix $S_{K}$ after K1K-1 iterations.
While this method confirms path existence, it treats a high-risk path (e.g., via SSH) as identical to a low-risk one (e.g., via a database port), which our probabilistic model addresses. Let $S_{1}=I\lor A$ and $S_{k+1}=S_{k}\lor(S_{k}A)$; write $S_{K}$ for the $K$-hop boolean reachability matrix.
3.3 Computing Compromise Probability (Probabilistic Method)
Our model calculates the $K$-hop compromise probability matrix, $P_{K}$, where $(P_{K})_{ij}$ is the total probability that an attacker starting at asset $i$ can compromise asset $j$ in at most $K$ hops. We define a specialized matrix exponentiation using two custom operators.
  1. 1.
    The Probabilistic Composition Product ($\otimes$): This operator, replacing matrix multiplication, calculates the probability of new compromise paths forming by propagating existing compromise probabilities through the pivot potentials in $W$. For matrices $X$ and $Y$, the product $Z=X\otimes Y$ is:
    $Z_{ij}=(X\otimes Y)_{ij}=1-\prod_{m=1}^{n}(1-X_{im}\times Y_{mj})$ (1)
  2. 2.
    The Probabilistic Union Operator ($\lor_{p}$): This operator, replacing matrix addition, correctly combines probabilities from independent sets of paths using the union rule. For matrices $X$ and $Y$, the union $Z=X\lor_{p}Y$ is:
    $Z_{ij}=(X\lor_{p}Y)_{ij}=1-(1-X_{ij})(1-Y_{ij})$ (2)
Stopping rule. Iterate until $\|P_{k+1}-P_{k}\|_{\infty}<\varepsilon$ or $k=K$. Cost. Each iteration is $O(n^{3})$ and GPU-friendly (dense ops).
Independence Assumption
It is important to note that the probabilistic composition operator ($\otimes$) implicitly treats paths through different intermediate assets ($m$) as probabilistically independent events. This is a standard assumption in many network models that makes the problem computationally tractable. We acknowledge that in some real-world scenarios, such as an attack using a single stolen credential, these paths may not be fully independent.
Conservative bound. For correlated routes, we can replace $1-\prod_{m}(1-X_{im}Y_{mj})$ with $\max_{m}(X_{im}Y_{mj})$ (lower bound).
Using these operators, we compute $P_{K}$ via the following iterative algorithm:
Step 1: Initialization (Base Case)
The 1-hop compromise probability matrix, $P_{1}$, is the adjacency matrix $A$ with a diagonal of 1s, as an asset that is the starting point of an attack is by definition compromised.
Step 2: Iteration
For each hop from $k=1$ to K1K-1, we calculate the next state, $P_{k+1}$. The new state is the probabilistic union of the current state, $P_{k}$, with all the new compromise paths that can be formed by one additional hop.
$P_{k+1}=P_{k}\lor_{p}(P_{k}\otimes W)$
Step 3: Final Result
After K1K-1 iterations, the resulting matrix is the final K-hop Compromise Probability Matrix, $P_{K}$.
Hop-1 semantics. We treat the first hop as pure network reachability: $P_{1}=I\lor_{p}A$. Pivot potentials in $W$ affect only subsequent hops ($k\geq 2$) via $P_{k}\otimes W$.
4.  Quantifying Susceptibility and Blast Radius
The final K-hop Compromise Probability Matrix, $P_{K}$, provides a detailed, asset-to-asset view of lateral movement exposure. To make this data actionable, we derive two high-level metrics: Lateral-Movement Susceptibility (LMSK) and Blast-Radius Estimate (BREK).
4.1 Lateral-Movement Susceptibility (LMSK)
Concept: This metric quantifies the overall ease of movement within the network. It answers the question: ”If we randomly select two different assets, what is the average probability that a successful lateral movement event can occur between them within $K$ hops?”
Formal Definition: The Lateral-Movement Susceptibility is the mean of all non-diagonal entries in the final compromise probability matrix, $P_{K}$. This focus on non-diagonal elements ensures the metric measures the probability of movement *between* distinct assets.
$\text{LMS}_{K}=\frac{1}{n(n-1)}\sum_{i\neq j}(P_{K})_{ij}$ (3)
where $n$ is the total number of assets.
Interpretation: LMSK is a score between 0 and 1. A score near 0 indicates a well-segmented network where lateral movement is highly improbable. A score near 1 indicates a ”flat” network where assets are highly entangled and movement is easy. This metric is ideal for tracking the network’s security posture over time and measuring the effectiveness of segmentation controls.
4.2 Blast-Radius Estimate (BREK)
Concept: This metric quantifies the expected scope of compromise from an attack. It answers the business-critical question: ”Assuming an attacker can start at any asset, what percentage of our entire network do we expect to be compromised?”
Formal Definition: The Blast-Radius Estimate is the mean of *all* entries in the final compromise probability matrix, $P_{K}$, expressed as a percentage. By including the diagonal elements (which are always 1), this metric correctly incorporates the initially compromised asset as part of the total compromise.
$\text{BRE}_{K}=\left(\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}(P_{K})_{ij}\right)\times 100\%$ (4)
Interpretation: BREK provides a direct estimate of the potential scope of a breach. A BREK of 30% signifies that, in an average-case attack scenario, 30% of the total network assets are expected to be compromised. This metric is designed to communicate exposure in clear, non-technical terms, helping to justify security investments and prioritize hardening efforts on the assets that contribute most to the overall blast radius.
5.  Formal Properties of the Model
Overview. We prove monotonicity, boundedness, and convergence; derive a $K$-hop series upper bound and a single-path product lower bound; recover boolean reachability in the unweighted limit; characterize the limit as the least fixed point $P^{*}=F(P)=P\oplus(P\otimes W)$; and show all results persist under service-class lifting.
5.1 Monotonicity
The iterative process is monotonically non-decreasing. For any step $k\geq 1$, every element in the probability matrix $P_{k+1}$ is greater than or equal to its corresponding element in $P_{k}$.
$\forall i,j:(P_{k+1})_{ij}\geq(P_{k})_{ij}$
This is guaranteed by the definition of the probabilistic union operator ($\lor_{p}$). This property ensures the model behaves logically, as the total probability of compromise only ever increases or remains constant as more paths are considered.
5.2 Boundedness
The values in the compromise probability matrix are always bounded within the interval $[0,1]$. The pivot potentials $\pi(s)$ are defined in $[0,1]$, and both the probabilistic composition ($\otimes$) and union ($\lor_{p}$) operators are closed over this interval. This property guarantees the stability of the model, ensuring that the outputs are always valid probabilities.
5.3 Convergence
The sequence of matrices $P_{1},P_{2},\dots$ is guaranteed to converge to a fixed-point matrix $P^{*}$, where further iterations no longer change the result. This is a direct consequence of the Monotonicity and Boundedness properties (a monotonically increasing sequence that is bounded from above must converge).
Unlike unweighted reachability, which converges in at most n1n-1 steps, the probabilistic model may require more iterations for the probability values to become fully stable. This is because paths longer than n1n-1 can still contribute a small but non-zero probability to the total, incrementally refining the final result. In practice, the algorithm is governed by a stopping rule: iterate until the change between successive matrices falls below a predefined tolerance threshold, $\varepsilon$, or a maximum number of hops $K$ is reached. Formally, the iteration stops when $\|P_{k+1}-P_{k}\|_{\infty}<\varepsilon$ or $k=K$.
By continuity of $\oplus$ and $\otimes$ on $[0,1]$, the limit $P^{*}$ satisfies $P^{*}=F(P^{*})$ with $F(P)=P\oplus(P\otimes W)$; thus $P^{*}$ is the least fixed point above $P_{1}$.
5.4 Refinement of Binary Reachability
This is the most critical property of our model. It establishes a formal link to the unweighted model and demonstrates the value of the probabilistic approach.
Let $S_{K}$ be the final reachability matrix from the unweighted (Boolean) method, and $P_{K}$ be the final matrix from our probabilistic method. The following soundness relationship holds:
$(P_{K})_{ij}>0\;\Rightarrow\;(S_{K})_{ij}=1$
Consequence (off-diagonal). For all $i\neq j$,
$0\leq(P_{K})_{ij}\leq(S_{K})_{ij}.$
Thus, Boolean reachability upper-bounds connectivity, while our probabilistic model provides a probability-weighted refinement (typically strictly smaller).
However, the unweighted model is a binary, all-or-nothing assessment. All reachable nodes are treated as equal. Our probabilistic model provides a continuous refinement of this binary view. It assigns a meaningful probability to each path, allowing us to differentiate between:
  • High-probability paths: A small number of connections that represent a clear and present danger.
  • Low-probability paths: A large number of theoretical connections that are practically irrelevant for an attacker.
This property of probabilistic refinement is what allows the model to effectively ”reduce” the set of relevant reachable nodes. By focusing only on paths above a meaningful probability threshold (e.g., $P_{ij}>0.1$), a security team can prioritize actions and ignore the noise of low-probability connections that would dominate a standard reachability analysis. This directly demonstrates the model’s utility in reducing complexity and focusing on what matters.
5.5 Connection to Algebraic Path Problems
Our construction follows the algebraic-path template over $[0,1]$ using the probabilistic-sum (noisy-OR) as the “addition” $\oplus$ and ordinary multiplication as the “multiplication” $\otimes$:
xy= $x\oplus y\;=\;1-(1-x)(1-y),\qquad x\otimes y\;$=\;xy.
Matrix composition then uses these operations in the usual way:
$(X\otimes Y)_{ij}\;=\;\bigoplus_{m=1}^{n}X_{im}\otimes Y_{mj}\;=\;1-\prod_{m=1}^{n}\bigl{(}1-X_{im}Y_{mj}\bigr{)},$
which is exactly our probabilistic composition operator. (Here $\oplus$ corresponds to the $\lor_{p}$ operator used elsewhere.)
Remark (semiring-like, non-distributive).
The pair $(\oplus,\otimes)$ is associative, commutative, and monotone with identities 0 (for $\oplus$) and $1$ (for $\otimes$), but it is not distributive. For all $a,b,c\in[0,1]$,
$(a\otimes b)\,\oplus\,(a\otimes c)\ \geq\ a\otimes(b\oplus c),\qquad(a\otimes c)\,\oplus\,(b\otimes c)\ \geq\ (a\oplus b)\otimes c,$
with equality iff, respectively, $a\in\{0,1\}$ or $bc=0$, and $c\in\{0,1\}$ or $ab=0$. We do not rely on distributivity in our results.
Fixed-point view.
Let $F(P)=P\oplus(P\otimes W)$. Since $F$ is monotone on the complete lattice $0,1]^{n\times n}$ and bounded in $[0,1]$, iterating
$P_{k+1}\;=\;P_{k}\oplus(P_{k}\otimes W)$
computes the path-closure (least fixed point) up to either a prescribed hop limit $K$ or an $\varepsilon$-tolerance, grounding our iteration in standard algebraic-path theory.
5.6 K-hop Series Bound for Lateral Propagation
Define, for $K\geq 1$, the path-sum matrix
$U_{K}\;=\;A\;+\;AW\;+\;AW^{2}\;+\;\cdots\;+\;AW^{K-1},$
where products are the usual matrix products. The term $AW^{t}$ aggregates all paths of length t+1t{+}1 (one landing hop in $A$ followed by $t$ pivot steps in $W$).
Proposition 1 (K-hop series upper bound).
For all $K\geq 1$ and all $i\neq j$,
$(P_{K})_{ij}\;\leq\;\min\!\bigl{(}1,\ (U_{K})_{ij}\bigr{)},\qquad\text{and}\qquad(P_{K})_{ii}=1.$
Proof. Fix $i\neq j$. Let $\mathcal{P}_{\leq K}(i\!\to\!j)$ be the set of all directed paths from $i$ to $j$ of length at most $K$. By construction, $(P_{K})_{ij}$ is the probability of the union of the path events $\{E_{p}\}_{p\in\mathcal{P}_{\leq K}}$. By Boole’s union inequality, $\Pr(\bigcup_{p}E_{p})\leq\sum_{p}\Pr(E_{p})$. For a path of length $L\geq 1$ with vertices $i\to m_{1}\to\cdots\to m_{L-1}\to j$, the model assigns probability $A_{im_{1}}\prod_{r=1}^{L-1}W_{m_{r}m_{r+1}}$, so summing over all such paths of fixed length $L$ yields $(AW^{L-1})_{ij}$. Summing over $L=1,\dots,K$ gives $(U_{K})_{ij}$. Capping at $1$ preserves validity. The diagonal equality $(P_{K})_{ii}=1$ follows from the initialization $P_{1}=I\lor_{p}A$. $\square$
Corollary 1 (infinite-hops bound).
If the spectral radius $\rho(W)<1$, then the Neumann series converges and
$U_{\infty}\;:=\;\sum_{t=0}^{\infty}AW^{t}\;=\;A(I-W)^{-1},\qquad\text{hence}\qquad(P_{\infty})_{ij}\;\leq\;\min\!\bigl{(}1,\ (U_{\infty})_{ij}\bigr{)}.$
Proof. Since $\rho(W)<1$, $\sum_{t\geq 0}W^{t}=(I-W)^{-1}$ converges. Apply Proposition 1 with $K\to\infty$ and linearity of matrix multiplication. $\square$
Proposition 2 (single-path lower bound).
Fix $K\geq L\geq 1$. Suppose there exists a directed path $p:\ i\!=\!m_{0}\to m_{1}\to\cdots\to m_{L-1}\to m_{L}\!=\!j$ in $A$ of length $L$, and let the pivot weights on its L1L-1 pivot steps be $w_{r}:=W_{m_{r}m_{r+1}}\in[0,1]$ for $r=1,\dots,L-1$. Then
$(P_{K})_{ij}\;\geq\;\prod_{r=1}^{L-1}w_{r}.$
Proof. $P_{K}(i,j)$ is the probability of a union over all $\leq K$-hop paths; it is at least the probability of any particular path event. Under the model’s hop-1 semantics ($A$ gives certainty on the first hop), the path $p$ succeeds with probability $\prod_{r=1}^{L-1}w_{r}$. $\square$
Corollary 2 (uniform privileged pivots).
Under the conditions of Proposition 2, if every pivot step on $p$ satisfies $w_{r}\geq p$ for some $p\in[0,1]$, then
$(P_{K})_{ij}\;\geq\;p^{\,L-1}.$
Proof. By monotonicity of products, $\prod_{r=1}^{L-1}w_{r}\geq\prod_{r=1}^{L-1}p=p^{L-1}$; combine with Proposition 2. $\square$
5.7 Service-Class Lifting (Exact Land-then-Pivot Semantics)
Not all footholds are equal: landing via an interactive remote-service (e.g., RDP/SSH) typically enables immediate lateral movement, whereas landing via application protocols (e.g., SQL/HTTP) rarely does without an exploit. To encode this service memory exactly, we lift states from nodes to (node, class) pairs.
Classes and per-class adjacency.
[…]
$A\;=\;\bigvee_{c\in\mathcal{C}}A^{(c)},\quad\text{with }A_{ii}=0.$
Let $\pi_{\mathrm{pivot}}(c)\in[0,1]$ denote the pivot probability conferred by a foothold acquired via class $c$.
Lifted transition.
A lifted state is $(j,c)$ meaning “at node $j$ with foothold class cc.” From $(j,c)$ one can move to $(\ell,c^{\prime})$ if and only if $A^{(c^{\prime})}_{j\ell}=1$, and the step succeeds with probability $\pi_{\mathrm{pivot}}(c)$. This defines the class-conditioned transition matrix $T\in[0,1]^{(n|\mathcal{C}|)\times(n|\mathcal{C}|)}$ with blocks
$T_{(j,c)\to(\ell,c^{\prime})}\;=\;\pi_{\mathrm{pivot}}(c)\,\mathbf{1}\{A^{(c^{\prime})}_{j\ell}=1\}.$
Initialization (hop 1: pure reach).
From a start node $i$, the first hop ignores pivot and only records the landing class:
$P^{\mathrm{lift}}_{1}\big{(}i\to(j,c)\big{)}\;=\;A^{(c)}_{ij},\qquad P^{\mathrm{node}}_{1}(i\to i)=1.$
Iteration (hops $\geq 2$).
For $k\geq 1$, propagate in the lifted space using the same operators as in Section 3.3:
$P^{\mathrm{lift}}_{k+1}\;=\;P^{\mathrm{lift}}_{k}\;\lor_{p}\;\big{(}P^{\mathrm{lift}}_{k}\;\otimes\;T\big{)}.$
Aggregation back to nodes.
The $K$-hop node-to-node compromise probability is obtained by class-wise probabilistic union at the destination:
$(P_{K})_{i\ell}\;=\;\bigl{(}\;\lor_{p}\;\{\,P^{\mathrm{lift}}_{K}(i\to(\ell,c))\ :\ c\in\mathcal{C}\,\}\bigr{)}.$
All metrics (LMSK, BREK) continue to use $P_{K}$ exactly as defined here.
Compatibility with the base model.
If classes are collapsed by setting $W_{j\ell}=1-\prod_{c\in\mathcal{C}}\bigl{(}1-\pi_{\mathrm{pivot}}(c)\,\mathbf{1}\{A^{(c)}_{j\ell}=1\}\bigr{)}$, the lifted model reduces to the node-level model. Conversely, the lifted construction preserves hop-1 semantics and makes the dependence on the landing class explicit at every subsequent hop.
Theorems unchanged.
Proposition 1 (K-hop series upper bound) and Proposition 2 (single-path lower bound), and their corollaries, hold verbatim under this lifting. In Proposition 2 the per-step weights become the class-conditioned values $w_{r}=\pi_{\mathrm{pivot}}(c_{r})$ determined by the foothold class present at the source of step $r$. In Proposition 1, the path-sum bound is obtained by summing over class-typed paths; the resulting bound is still computable via ordinary matrix products by replacing $A$ with the block structure induced by $\{A^{(c)}\}$ and $W$ with $T$.
5.8 Computational Cost
Each iteration of the algorithm is dominated by the probabilistic composition product ($\otimes$), which has a computational complexity of $O(n^{3})$, equivalent to standard dense matrix multiplication. The operations involved (primarily multiplications and additions) are highly parallelizable and operate on dense matrices. This makes the algorithm well-suited for hardware acceleration on platforms such as GPUs, ensuring its practical applicability to real-world networks of significant size.
6.  Examples and Evidence
In this section, we provide a practical demonstration of our model. We begin with a simple, illustrative example to clarify the computational steps. We then present a more complex scenario to provide evidence of the model’s utility in identifying non-obvious security risks that are missed by standard reachability analysis.
6.1 Illustrative Example
Consider a simple network of three assets: a Web Server (Asset 1), an Application Server (Asset 2), and a Database (Asset 3). For this example, the Adjacency Matrix ($A$) representing the reachable services, and the corresponding Pivot Potential Matrix ($W$) derived from those services, are defined as:
$A=\begin{bmatrix}0&1&0\\ 1&0&1\\ 0&0&0\end{bmatrix}\quad W=\begin{bmatrix}0.0&0.2&0.0\\ 0.9&0.0&0.3\\ 0.0&0.0&0.0\end{bmatrix}$
Using the iterative algorithm defined in Section 3, we compute the 2-hop compromise probability matrix, $P_{2}$:
  • First, the 1-hop matrix $P_{1}$ is initialized from $A$.
  • Then, we compute the 2-hop result: $P_{2}=P_{1}\lor_{p}(P_{1}\otimes W)$.
This yields the final probability matrix:
$P_{2}=\begin{bmatrix}1.0&1.0&0.3\\ 1.0&1.0&1.0\\ 0.0&0.0&1.0\end{bmatrix}$
From this matrix, we calculate the final metrics for $K=2$:
  • LMS2: The average of the 6 off-diagonal elements is $(1.0+0.3+1.0+1.0+0.0+0.0)/6=0.55$.
  • BRE2: The average of all 9 elements is $(6.3)/9\times 100\%=70\%$.
This simple example demonstrates how the high pivot potential of the connection from Asset 2 to 1 creates a high probability of compromise across the network, resulting in a high Blast-Radius Estimate.
6.2 Evidence of Practical Utility
We now provide evidence of the model’s value in a more complex scenario where the goal is to prioritize security controls. Consider a network with four assets: a User Workstation (Asset 1), a Web Server (Asset 2), a critical Database (Asset 3), and a Domain Controller (Asset 4).
The reachable services are:
  • Workstation (1) $\to$ Web Server (2) via a low-privilege web port, $\pi=0.1$.
  • Web Server (2) $\to$ Database (3) via a standard SQL port, $\pi=0.2$.
  • Workstation (1) $\to$ Domain Controller (4) via an RDP port left open for remote admin access, $\pi=0.9$.
Analysis with the Unweighted Method: A standard reachability analysis (using the Boolean method) would identify two lateral movement paths originating from the Workstation: one to the Database (length 2) and one to the Domain Controller (length 1). Both are considered ”reachable.” A security team might mistakenly prioritize the longer path involving the Web Server, or treat both paths as equally important, failing to see the true concentration of exposure.
Analysis with Our Probabilistic Model: Our model calculates the compromise probabilities. The 2-hop compromise matrix $P_{2}$ would show:
  • The probability of compromising the Database from the Workstation, $(P_{2})_{13}$, would be low. The probability of the two-hop path is based on the composition of low-privilege pivots (approximately $0.1\times 0.2=0.02$).
  • The probability of compromising the Domain Controller, $(P_{2})_{14}$, would be very high, dominated by the single high-privilege RDP hop (0.9).
The BRE from the Workstation would be heavily weighted by the near-certain compromise of the Domain Controller.
6.3 Illustrative Field Example: Service-Class Contrast in an Enterprise Snapshot
We complement our formal results with an illustrative observation from a production environment described by Tyagi [2025]. The analysis compares adjacency induced by two services: RDP (TCP/3389), an interactive remote-admin channel, and Epic Cache (TCP/6966), an application-specific service. The induced RDP adjacency was reported as nearly complete (most node pairs reachable directly), whereas the Epic Cache adjacency was nearly empty (few pairs reachable). Under our framework, this service-class asymmetry implies substantially higher step-success and union mass for RDP-driven paths, elevating both LMSK (off-diagonal mean) and BREK (mean including the diagonal) relative to app-only services.
The same field note underscores that adversaries frequently achieve lateral movement using valid accounts over remote services (RDP/SSH/SMB/WinRM) without relying on unpatched CVEs, aligning with ATT&CK’s “Remote Services” technique and our CVE-agnostic design. We emphasize that this is a single-environment qualitative example (not a statistical study). Nevertheless, the observation is consistent with our model’s predictions and with practitioner guidance prioritizing constraints on interactive remote services. [Tyagi, 2025; MITRE, 2025b, f, c]
Conclusion: This scenario provides clear evidence of the model’s value. While the unweighted method presents a noisy picture of undifferentiated risks, our probabilistic model immediately and quantitatively identifies the single RDP connection as the most significant threat to the network. It correctly filters the noise from the low-probability path and pinpoints the connection that must be prioritized for remediation, demonstrating its practical utility in guiding security decisions.